# Recent Advances in Speech Language Models: A Survey

### v1

@@ 3.1.1 Semantic Understanding Objective
A semantic understanding speech tokenizer typically comprises a speech encoder and a quantizer, where the speech encoder encodes the essential information from the waveform and the quantizer discretizes continuous representations into discrete tokens. Let $f_{E}(\cdot)$ denote the speech encoder parameterized by $\theta_{f_{E}}$, we have $\textbf{v}=f_{E}(\textbf{a};\theta_{f_{E}})$, where $\textbf{v}=(v_{1},v_{2},\ldots,v_{P})$ represents the encoded representations. Since v is still continuous, a quantizer $d(\cdot)$ is utilized to discretize the representation. Depending on different design choices, the discrete speech tokens $\textbf{s}=(s_{1},s_{2},\ldots,s_{P})$ can either be derived from a or v. Therefore, we have $\textbf{s}=d(\textbf{v};\theta_{d})$ or $\textbf{s}=d(\textbf{a};\theta_{d})$. After that, s can be used to train the speech tokenizer as a target label (such as masking $_{\text{mask}}\subset\textbf{a}$ a and reconstructing its corresponding label $_{\text{mask}}\subset\textbf{s}$ s Hsu et al. (2021)) or to train the following language model.
@@
The key design choices lie in how to effectively encode and quantize the speech into discrete tokens. Wav2vec 2.0 Baevski et al. (2020) uses a convolutional encoder followed by a product quantization module Jegou et al. (2010) to discretize the continuous waveform. Then, a portion of the quantized representations is masked and modeled using a contrastive loss. W2v-BERT Chung et al. (2021) is built upon wav2vec 2.0 and proposes to use Masked Language Modeling (MLM) loss Devlin (2018) in addition to contrastive loss. Similarly, HuBERT Hsu et al. (2021) uses the k-means algorithm to cluster the speech utterances into a certain number of hidden units, and then perform MLM to predict the target hidden units from the masked speech utterances. To better align the representation of text and speech modalities, Google USM Zhang et al. (2023c) utilizes text-injection loss Chen et al. (2022b) at the second pre-training stage to improve the performance and robustness of the downstream tasks. WavLM Chen et al. (2022a) adds the speech denoising objective during pre-training. While the majority of speech tokenizer studies focus on semantic-related tasks such as ASR and TTS, WavLM shows that speech denoising can boost the performance of non-semantic tasks such as speaker verification and speech separation. A full list of downstream tasks is listed in section 5.
3.1.2 Acoustic Generation Objective
@@
Speech tokenizers with an acoustic generation objective focus on capturing the acoustic features necessary for generating high-quality speech waveforms. These tokenizers prioritize the preservation of essential acoustic characteristics over semantic content, making them suitable for speech synthesis tasks.
@@
To generate high-quality speech waveforms, acoustic generation speech tokenizers employ a speech synthesis or speech reconstruction objective. To achieve this, the architecture typically includes an encoder, a quantizer, and a decoder. Same as before, the encoder $f_{E}(\cdot)$ and quantizer $d(\cdot)$ transform the original waveform into discrete tokens. After that, the decoder $f_{D}(\cdot)$ reconstructs these tokens back into speech waveforms. This process is represented by $\hat{\textbf{a}}=f_{D}(\textbf{s};\theta_{f_{E}})$, where $\hat{\textbf{a}}$ is the generated or reconstructed waveform.
@@
Neural audio codecs are very suitable for and are primarily employed as acoustic generation speech tokenizers. These codecs utilize the advanced modeling capabilities of deep neural networks to compress audio signals into a compact representation, typically in the form of quantized tokens. For example, both SoundStream Zeghidour et al. (2021) and EnCodec Défossez et al. (2022) use convolution blocks as the encoder and use Residual Vector Quantization (RVQ) Zeghidour et al. (2021) as the quantizer. This mechanism allows for codecs to efficiently transmit or store audio with minimal loss in quality. Since the output of codecs is in discrete format, they can also be leveraged by SpeechLM to autoregressively generate speech.
3.1.3 Mixed Objective
@@
Speech tokenizers with a mixed objective aim to balance both semantic understanding and acoustic generation. Currently, the development of these tokenizers is in its early stages. Most existing mixed speech tokenizers primarily adopt the architecture of acoustic generation speech tokenizers and focus on distilling information from semantic tokenizers into the acoustic tokenizer. SpeechTokenizer Zhang et al. (2023b) utilizes the RVQ-GAN Défossez et al. (2022); Zeghidour et al. (2021) architecture, distilling semantic information from HuBERT Hsu et al. (2021) to the first layer of RVQ. Building on SpeechTokenizer, Mimi Défossez et al. (2024) employs a single vector quantizer (VQ) to extract information from WavLM Chen et al. (2022a) and incorporates another RVQ module to learn the acoustic information.
@@
Moshi Défossez et al. (2024) Mimi (SeaNet Tagliasacchi et al. (2020) + Residual Vector Quantizer Zeghidour et al. (2021)) Transformer* Mimi
VITA Fu et al. (2024) CNN + Transformer + MLP Fu et al. (2024) Mixtral Jiang et al. (2024) Text-to-Speech Toolkit Fu et al. (2024)
LSLM Ma et al. (2024) vq-wav2vec Baevski et al. (2019) Decoder-Only Transformer Vaswani (2017) UniVATS Du et al. (2024a)
SpiRit-LM Nguyen et al. (2024) HuBERT, VQ-VAE, speechprop LLaMA 2 Touvron et al. (2023b) HiFi-GAN Kong et al. (2020a); Polyak et al. (2021)
TWIST Hassid et al. (2024) HuBERT Hsu et al. (2021) OPT Zhang et al. (2022b), LLaMA Touvron et al. (2023a) HiFi-GAN
VOXTLM Maiti et al. (2024) HuBERT OPT Zhang et al. (2022b) HiFi-GAN
Voicebox Le et al. (2024) EnCodec Défossez et al. (2022) Transformer* Vaswani (2017) HiFi-GAN
VioLA Wang et al. (2024) EnCodec Transformer* Codec Decoder Défossez et al. (2022)
FunAudioLLM SpeechTeam (2024) SAN-M Gao et al. (2020) Transformer* HiFTNet Li et al. (2023)
AudioChatLlama Fathullah et al. (2024) Conformer layers Gulati et al. (2020) LLaMA 2 -
Wav2Prompt Deng et al. (2024) Conformer* LLaMA-2, Vicuna Chiang et al. (2023) -
SpeechVerse Das et al. (2024) WavLM Large Chen et al. (2022a), Best-RQ Chiu et al. (2022) Flan-T5-XL Chung et al. (2024) -
SpeechGPT-Gen Zhang et al. (2024) SpeechTokenizer Zhang et al. (2023b) LLaMA 2 SpeechTokenizer decoder Zhang et al. (2023b)
LauraGPT Du et al. (2023) Conformer* Qwen Bai et al. (2023) Transformer + Codec Decoder
Spectron Nachmani et al. (2023) Conformer* PaLM 2* Anil et al. (2023a) WaveFit Koizumi et al. (2023)
AudioLM Borsos et al. (2023) w2v-BERT Chung et al. (2021) Decoder-Only Transformer* SoundStream* Zeghidour et al. (2021)
@@
UniAudio Yang et al. (2023b) EnCodec, Hifi-codec Yang et al. (2023a), Improved RVQGAN Kumar et al. (2023) Transformer* Codec Decoder
VALL-E Wang et al. (2023a) EnCodec GPT-3 Brown et al. (2020) Codec Decoder
LTU Gong et al. (2023) Audio Spectrogram Transformer Gong et al. (2021) Vicuna -
Llama-Omni Fang et al. (2024) Whisper Radford et al. (2023) Llama-3.1 Dubey et al. (2024b) HiFi-GAN
SALMONN Tang et al. (2024) Whisper, BEATs Chen et al. (2023) Vicuna -
Mini-Omni Xie and Wu (2024) Whisper + ASR Adapter Xie and Wu (2024) Qwen2 Yang et al. (2024) TTS Adapter Xie and Wu (2024)
COSMIC Pan et al. (2023) Whisper LLaMA-2 -
SpeechGPT Zhang et al. (2023a) HuBERT LLaMA HiFi-GAN
DLM Nguyen et al. (2023b) HuBERT Dialogue Transformer Nguyen et al. (2023b) HiFi-GAN
SUTLM Chou et al. (2023) HuBERT Transformer* -
dGSLM Nguyen et al. (2023b) HuBERT Transformer* HiFi-GAN
pGSLM Kharitonov et al. (2021) HuBERT MS-TLM Kharitonov et al. (2021) HiFi-GAN
GSLM Lakhotia et al. (2021) HuBERT, CPC Oord et al. (2018), Wav2vec 2.0 Baevski et al. (2020) Transformer* Tacotron-2 + Waveglow Shen et al. (2018); Prenger et al. (2019)
Table 1: Summarization of the architectural choice of speech tokenizer, language model, and vocoder in popular SpeechLMs. “-” represents non-existence or not indicated, * means the architecture is mainly based on the written one, “A, B” means the authors experimented with both A and B as the component, and “A + B” means “A” and “B” are combined to serve as the component.
3.2 Language Model
@@
Due to the success of TextLMs Achiam et al. (2023); Team et al. (2023); Dubey et al. (2024a), most SpeechLMs follow their architectures. They primarily employ transformers Vaswani (2017) or decoder-only architectures (such as OPT Zhang et al. (2022b), LLaMA Touvron et al. (2023a)) to generate speech in an autoregressive manner. To formally define it, given $|V_{t}|$ as the vocabulary size and $h$ as the hidden dimension, a typical text-based decoder-only transformer language model consists of an embedding matrix $E_{t}\in\mathbb{R}^{|V_{t}|\times h}$, a sequence of $L$ transformer decoder blocks $\textbf{De}=\{De_{1},De_{2},\ldots,De_{L}\}$, and an output embedding matrix $E^{\prime}_{t}\in\mathbb{R}^{h\times|V_{t}|}$. Therefore, the language model (LM) can be represented as
@@
To adapt the language model to generate speech, the original text tokenizer is changed to the speech tokenizers illustrated in section 3.1. $E_{t}\in\mathbb{R}^{|V_{t}|\times h}$ is thus changed to a speech embedding matrix $E_{s}\in\mathbb{R}^{|V_{s}|\times h}$, where $|V_{s}|$ represents the vocabulary size of the speech tokenizer. The output embedding matrix is also changed from $E^{\prime}_{t}\in\mathbb{R}^{h\times|V_{t}|}$ to $E^{\prime}_{s}\in\mathbb{R}^{h\times|V_{s}|}$. As a result, the language model in an SpeechLM is represented as
@@
Because the language model architecture of SpeechLMs is borrowed from TextLMs, it is natural that the resulting model can jointly model both text and speech modalities Nguyen et al. (2024); Zhang et al. (2023a). To achieve this, a naive and most adopted approach is to expand the vocabulary of the original TextLM to incorporate both text and speech tokens. Specifically, the speech embedding matrix is usually appended to the end of the text embedding matrix, resulting in a larger embedding matrix $E_{m}\in\mathbb{R}^{(|V_{t}|+|V_{s}|)\times h}$. Let m be a token sequence containing both speech and text tokens, the resulting language model becomes
@@
By doing so, the model can generate both text and speech in a single sequence, enabling much more diverse applications (see section 5).
3.3 Token-to-Speech Synthesizer (Vocoder)
@@
$\textbf{a}=V(\textbf{s};\theta_{V}),$ (5)
where $V$ is the vocoder model parameterized by $\theta_{V}$.
The pipeline of the SpeechLM vocoder can vary depending on the underlying vocoder model. There are two main pipelines: Direct synthesis and input-enhanced synthesis. Direct synthesis is the pipeline where the vocoder directly converts speech tokens generated by the language model into audio waveforms. For example, Polyak et al. (2021) adapts the HiFi-GAN Kong et al. (2020a) architecture and takes discrete tokens as inputs. In contrast, input-enhanced synthesis employs an additional module to transform the tokens into a continuous latent representation before they are fed into the vocoder Anastassiou et al. (2024); Betker (2023). The main reason for using this pipeline is that vocoders typically require intermediate audio representations, such as mel-spectrograms Kumar et al. (2019); Kong et al. (2020a); Lee et al. (2022), as input. When comparing the two pipelines, direct synthesis is generally simpler and faster than Input-Enhanced Synthesis. However, the choice of pipeline depends on the type of tokens used as input. Tokens from acoustic generation tokenizers contain sufficient acoustic information, making them suitable for sirect aynthesis. Conversely, tokens from semantic understanding tokenizers provide rich semantic information but lack fine acoustic details, particularly in higher frequencies. Therefore, these tokens are better enhanced into an acoustic-rich representation, such as mel-spectrograms, before synthesizing the final speech.
@@
3.3.1 GAN-based Vocoder
Generative Adversarial Network (GAN) is the most adopted architecture of the vocoders Kumar et al. (2019); Kong et al. (2020a); Polyak et al. (2021); Kim et al. (2021); Lee et al. (2022). It is well known for its fast and high-fidelity generation in speech synthesis tasks. The architecture of GAN includes a generator and a discriminator. Specifically, the generator creates realistic audio waveforms from random noise or input features, while the discriminator evaluates the authenticity of the generated audio against real audio samples.
To utilize GAN to synthesize high-fidelity speech, various training objectives are designed, focusing on different aspects. First, GAN loss is utilized as the fundamental objective for the operation of the generator and the discriminator. Specifically, the typical choice GAN loss for the generator ($G$) and discriminator ($D$) is to use the least squares loss function. The GAN loss for the generator ($\mathcal{L}_{\text{GAN}}(G;D)$) and the discriminator ($\mathcal{L}_{\text{GAN}}(D;G)$) are

### v2

@@ III-A1 Semantic Understanding Objective
[…]
@@
The key design choices lie in how to effectively encode (and quantize) speech into tokens. Wav2vec 2.0 [17] uses a convolutional encoder followed by a product quantization module [57] to discretize the continuous waveform. Then, a portion of the quantized representations is masked and modeled using a contrastive loss. W2v-BERT [19] is built upon wav2vec 2.0 and proposes to use Masked Language Modeling (MLM) loss [58] in addition to contrastive loss. Similarly, HuBERT [20] uses the k-means algorithm to cluster the speech utterances into a number of hidden units, and then perform MLM to predict the target hidden units from the masked speech utterances. To better align the representation of text and speech modalities, Google USM [23] utilizes text-injection loss [59] at the second pre-training stage to improve the performance and robustness of the downstream tasks. WavLM [24] adds the speech denoising objective during pre-training. While the majority of speech tokenizer studies focus on semantic-related tasks such as ASR and TTS, WavLM shows that speech denoising can boost the performance of non-semantic tasks such as speaker verification and speech separation. Section V presents a full list of downstream tasks.
III-A2 Acoustic Generation Objective
@@
Speech tokenizers with an acoustic generation objective focus on capturing the acoustic features necessary for generating high-quality speech waveforms. These tokenizers prioritize the preservation of essential acoustic characteristics over semantic content, making them suitable for speech (re)synthesis tasks.
@@
To generate high-quality speech waveforms, acoustic generation speech tokenizers employ a speech synthesis or speech reconstruction objective. To achieve this, the architecture typically includes an encoder, a quantizer, and a decoder. Same as before, the encoder $f_{E}(\cdot)$ and quantizer $d(\cdot)$ transform the original waveform into tokens. After that, the decoder $f_{D}(\cdot)$ reconstructs these tokens back into speech waveforms. This process is represented by $\hat{\textbf{a}}=f_{D}(\textbf{s};\theta_{f_{E}})$, where $\hat{\textbf{a}}$ is the generated or reconstructed waveform.
@@
Neural audio codecs are very suitable for and are primarily employed as acoustic generation speech tokenizers [22, 36]. These codecs utilize the advanced modeling capabilities of deep neural networks to compress audio waveforms into a compact representation, typically in the form of discrete tokens. Using the encoder-quantizer-decoder architecture, the encoder compresses the audio into latent representations, the quantizer discretizes these representations (commonly through vector quantization (VQ) [60] or residual vector quantization (RVQ) [22]), and the decoder reconstructs the discrete tokens back into audio waveforms. Therefore, the encoder and/or the quantizer are utilized as an acoustic speech tokenizer.
III-A3 Mixed Objective
@@
Speech tokenizers with a mixed objective aim to balance both semantic understanding and acoustic generation. The goal is to harness the advantages of both types of tokenizers. Currently, the development of these tokenizers is in its early stages. Most existing mixed speech tokenizers primarily adopt the architecture of acoustic generation speech tokenizers and focus on distilling information from semantic tokenizers into the acoustic tokenizer. SpeechTokenizer [21] utilizes the RVQ-GAN [36, 22] architecture, distilling semantic information from HuBERT [20] to the first layer of RVQ. Inspired by SpeechTokenizer, Mimi [9] employs a single VQ to extract information from WavLM [24] and incorporates another RVQ module to learn the acoustic information.
TABLE II: Summarization of the architectural choice of speech tokenizer, language model, and vocoder in popular SpeechLMs. “-” represents non-existence or not indicated, * means the architecture is mainly based on the written one, “A, B” means the authors experimented with both A and B as the component, and “A + B” means “A” and “B” are combined to serve as the component.
@@
SLAM-Omni [41] Whisper Encoder  [12] + Linear Projector Qwen2 [28] -
OmniFlatten [40] CosyVoice Encoder [61] Qwen2 CosyVoice Decoder
SyncLLM [62] HuBERT [20] LLaMA-3 [2] HiFi-GAN [34, 35]
EMOVA [63] SPIRAL [64] LLaMA-3 VITS [65]
Freeze-Omni [53] Transformer [25] Qwen2 TiCodec [66]
IntrinsicVoice [67] HuBERT Qwen2 HiFi-GAN
Mini-Omni2 [52] Whisper Qwen2 Mini-Omni [10]
SALMONN-omni [56] Mamba Streaming Encoder [68] - VoiceCraft [69] + Codec Decoder
Zeng et al. [70] Whisper + ASR GLM [29] CosyVoice
Parrot [45] VQ-VAE LLaMA-3, Mistral, Gemma 2 HiFi-GAN
GPST [71] EnCodec [36] Transformer Codec Decoder
Moshi [9] Mimi [9] Transformer* Mimi
VITA [55] CNN + Transformer + MLP [55] Mixtral [30] Text-to-Speech Toolkit [55]
LSLM [51] vq-wav2vec [18] Decoder-Only Transformer UniVATS [72]
SpiRit-LM [5] HuBERT, VQ-VAE [60], speechprop LLaMA-2 [27] HiFi-GAN
TWIST [38] HuBERT OPT [3], LLaMA [26] HiFi-GAN
VOXTLM [73] HuBERT OPT [3] HiFi-GAN
Voicebox [74] EnCodec Transformer* [25] HiFi-GAN
VioLA [43] EnCodec Transformer* Codec Decoder [36]
FunAudioLLM [75] SAN-M [76] Transformer* HiFTNet [77]
SpeechGPT-Gen [46] SpeechTokenizer [21] LLaMA-2 SpeechTokenizer decoder [21]
LauraGPT [49] Conformer* Qwen [78] Transformer + Codec Decoder
Spectron [47] Conformer* PaLM 2* [79] WaveFit [80]
AudioLM [81] w2v-BERT [19] Decoder-Only Transformer* SoundStream* [22]
@@
UniAudio [82] EnCodec, Hifi-codec [83], Improved RVQGAN [84] Transformer* Codec Decoder
Llama-Omni [11] Whisper LLaMA-3.1 HiFi-GAN
Mini-Omni [10] Whisper + ASR Adapter [10] Qwen2 TTS Adapter [10]
tGSLM [48] Segmentation + SSE [85] + Lexical embedder Transformer* Tacotron-2 + Waveglow [32, 33]
SpeechGPT [8] HuBERT LLaMA HiFi-GAN
dGSLM [4] HuBERT Dialogue Transformer [4] HiFi-GAN
SUTLM [50] HuBERT Transformer* -
pGSLM [42] HuBERT MS-TLM [42] HiFi-GAN
GSLM [37] HuBERT, CPC [16], Wav2vec 2.0 [17] Transformer* Tacotron-2 + Waveglow
III-B Language Model
@@
Due to the success of TextLMs [1, 86, 2], most SpeechLMs follow their architectures. They primarily employ transformers [25] or decoder-only architectures (such as OPT [3], LLaMA [26]) to generate speech in an autoregressive manner. To formally define it, given $|V_{t}|$ as the vocabulary size and $h$ as the hidden dimension, a typical text-based decoder-only transformer language model consists of an embedding matrix $E_{t}\in\mathbb{R}^{|V_{t}|\times h}$, a sequence of $L$ transformer decoder blocks $\textbf{De}=\{De_{1},De_{2},\ldots,De_{L}\}$, and an output embedding matrix $E^{\prime}_{t}\in\mathbb{R}^{h\times|V_{t}|}$. Therefore, the language model (LM) can be represented as
@@
To adapt the language model to generate speech, the original text tokenizer is changed to the speech tokenizers illustrated in section III-A. When using discrete tokens, $E_{t}\in\mathbb{R}^{|V_{t}|\times h}$ is changed to a speech embedding matrix $E_{s}\in\mathbb{R}^{|V_{s}|\times h}$, where $|V_{s}|$ represents the vocabulary size of the speech tokenizer. The output embedding matrix is also changed from $E^{\prime}_{t}\in\mathbb{R}^{h\times|V_{t}|}$ to $E^{\prime}_{s}\in\mathbb{R}^{h\times|V_{s}|}$. As a result, the language model in a SpeechLM is represented as
@@
Because the language model architecture of SpeechLMs is borrowed from TextLMs, it is natural that the resulting model can jointly model both text and speech modalities [5, 8]. To achieve this, a naive and most adopted approach is to expand the vocabulary of the original TextLM to incorporate both text and speech tokens. Specifically, the speech embedding matrix is usually appended to the end of the text embedding matrix, resulting in a larger embedding matrix $E_{m}\in\mathbb{R}^{(|V_{t}|+|V_{s}|)\times h}$. Let m be a token sequence containing both speech and text tokens, the resulting language model becomes
@@
By doing so, the model can generate both text and speech in a single sequence, enabling much more diverse applications (see section V). In contrast, when modeling with continuous tokens, the embeddings derived from the speech tokenizer are directly fed into the language model. In this case, the architecture of the language model remains unchanged.
III-C Token-to-Speech Synthesizer (Vocoder)
@@
$\textbf{a}=Vo(\textbf{s};\theta_{V_{o}}),$ (5)
where $Vo$ is the vocoder model parameterized by $\theta_{V_{o}}$.
The pipeline of the SpeechLM vocoder can vary depending on the underlying vocoder model. There are two main pipelines: Direct synthesis and input-enhanced synthesis. Direct synthesis is the pipeline where the vocoder directly converts speech tokens generated by the language model into audio waveforms. For example, Polyak et al. [35] adapts the HiFi-GAN [34] architecture and takes speech tokens as inputs. In contrast, input-enhanced synthesis employs an additional module to transform the tokens into a continuous latent representation before they are fed into the vocoder [87, 88]. The main reason for using this pipeline is that vocoders typically require intermediate audio representations, such as mel-spectrograms [89, 34, 90], as input. For example, CosyVoice [61] introduces a Conditional Flow-Matching (CFM) model to convert speech tokens into mel-spectrogram, and then leverages a HiFi-GAN to synthesize the final waveform. When comparing the two pipelines, direct synthesis is generally simpler and faster than input-enhanced synthesis. However, the choice of pipeline depends on the type of tokens used as input. Tokens from acoustic generation tokenizers contain sufficient acoustic information, making them suitable for direct synthesis. Conversely, tokens from semantic understanding tokenizers provide rich semantic information but lack fine acoustic details, particularly in higher frequencies. Therefore, these tokens are better enhanced into an acoustic-rich representation, such as mel-spectrograms, before synthesizing the final speech.
@@
III-C1 GAN-based Vocoder
Generative Adversarial Network (GAN) is the most adopted architecture of the vocoders [89, 34, 35, 91, 90]. It is well known for its fast and high-fidelity generation in speech synthesis tasks. The architecture of GAN includes a generator and a discriminator. Specifically, the generator creates realistic audio waveforms from random noise or input features, while the discriminator evaluates the authenticity of the generated audio against real audio samples.
To utilize GAN to synthesize high-fidelity speech, various training objectives are designed, focusing on different aspects. First, GAN loss is utilized as the fundamental objective for the operation of the generator and the discriminator. Specifically, the typical choice of GAN loss for the generator ($G$) and discriminator ($D$) is to use the least squares loss function. The GAN loss for the generator ($\mathcal{L}_{\text{GAN}}(G;D)$) and the discriminator ($\mathcal{L}_{\text{GAN}}(D;G)$) are